• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Armstrong

    Log entry time 23:57:11 on July27,2003

    Entry number 107563

    keyword=Target Shift Summary


    Target Log   Swing shift   July 27 2003



    D.S. Armstrong




    Shift Start:

    holding field: 0°
    target HWP: out
    lasers: 4,6,7 on
    Target: Pol 3He
    Cell: Priapus
    Beam: 4.7 µ A





       17:13   ask for beam off
       17:18   turn off lasers 4,6,7
       17:20   start rotation 0° to 270°.
       17:23   turn on Transverse lasers 1,2,3
       17:26   ask for beam back

       17:53   ask for beam off
       17:55   Turn off Transverse lasers (1,2,3)
       17:57-18:00   rotate 270° to 0°
       18:01   Turn on Laser 4
       18:02   Turn on Laser 6
       18:03   Turn on Laser 7

       18:04-18:10   Move to reference cell (empty)

       18:48-18:58   Empty Reference cell, Run 2913

       18:59   fill Ref. cell with nitrogen to 140 psig

       19:07-19:17   Full Reference cell, Run 2914

    can't move target yet, since Compton people are using the beam, so
    take another reference cell run in the meantime.

       19:40   Compton done for now; start move to
    3He cell

       19:41   Vent and evacuate Reference cell while target
    ladder moving

       19:45   At 3He cell position;
    ask for 4.7 µA beam.

       20:56 &nbp; ask for no beam, to rotate to 270°

       20:57   Turn off lasers 4,6,7

       20:58   Start rotating 0° to 270 °

       21:03   rotation complete, turn on Laser 1

       21:04   turn on Laser 2

       21:05   turn on Laser 3

       21:05   ask for beam back

       21:33   ask for beam off, to go to NMR measurement and make
    kinematic change

       21:34   Transverse Lasers off

       21:35   rotate holding field 270 ° to 0 °

          wait for temperature to stabilize before NMR


       22:00   make NMR measurement

       22:02   start moving to Reference cell (empty) position

       22:05   At empty Reference cell position; ask for 5 µA so Compton

    can run (we are waiting for dipole)

          NMR result = 40.05 %
    polarization


       22:10   Dipole is set; start run with 10 µA on Empty Reference
    cell.

       22:25   Fill reference cell to 140 psig

       22:27   Start run at 5.3 µA, with Nitrogen Reference Cell

       22:37   ask for beam off to move to 3He cell.

        (aside: it seems to take several minutes between the time
    the beam goes off and when the target motion FSD mask comes on; have to
    call MCC again to get FSD mask on)


       22:40   start moving to 3He cell position; while target
    is moving, vent and evacuate Reference cell.

       22:45   3He cell in position, ask for 5 µA

       23:28   Ask for beam off to rotate field

       23:29   Turn off Longitudinal lasers (4,6,7)

       23:30-23:33   Rotate holding field 0° to 270 °

       23:34   Turn on Laser 1

       23:35   Turn on Laser 2

       23:36   Turn on Laser 3

    almost finished the 270 ° setting, when beam off (FSD trip).




    Shift End:

    holding field: 270 °
    target HWP: out
    lasers: transverse: 1,2,3 ON
    Target: Pol 3He
    Cell: Priapus
    Beam: off (FSD trip)



    hand off to owl shift, go home to sleep...